# Universal Quantum Tomography With Deep Neural Networks

### v2

@@ Abstract
Quantum state tomography is a crucial technique for characterizing the state of a quantum system, which is essential for many applications in quantum technologies. In recent years, there has been growing interest in leveraging neural networks to enhance the efficiency and accuracy of quantum state tomography. Still, many of them did not include mixed quantum state, since pure states are arguably less common in practical situations. In this research paper, we present two neural networks based approach for both pure and mixed quantum state tomography: Restricted Feature Based Neural Network and Mixed States Conditional Generative Adversarial Network, evaluate its effectiveness in comparison to existing neural based methods. We demonstrate that our proposed methods can achieve state-of-the-art results in reconstructing mixed quantum states from experimental data. Our work highlights the potential of neural networks in revolutionizing quantum state tomography and facilitating the development of quantum technologies.
@@ I Introduction
Quantum State Tomography (QST) [15, 16, 17] aims to ascertain the unknown quantum state by conducting measurements on a finite set of identical copies of the system. If the state is characterized by the density matrix $\varrho$, residing in a $d$-dimensional Hilbert space, approximately $O(d/\varepsilon)$ copies are necessary to achieve an estimate of $\varrho$ with an error (measured as total variation distance) less than $\varepsilon$[18]. This underscores the considerable resource demands of QST for large-scale systems.
In a broader context, QST can be viewed as an inverse problem[19, 20]. Consequently, linear inversion[21] is often considered the most straightforward approach. However, it has drawbacks, such as the potential to yield a non-physical state and the inability to analytically determine the mean squared error bound of the estimate. To address these limitations, several alternative QST methods have been developed, including maximum-likelihood estimation (MLE)[22, 23], Bayesian tomography[24, 25], compressed sensing techniques[26, 27], and matrix-product states[28, 29], where MLE remains the most commonly employed approach.
@@
In our research, we introduce QST with 2 type of neural networks (NNs) architecture: Restricted Feature Based Neural Network (RFB-Net) and Mixed States Conditional Generative Adversarial Network (MS-CGAN). RFB-Net is a modified QST-NN[38] with modified regression modules that can classify quantum states into specific specific class and pointing out its features of the quantum state within certain constraints, finally using these features and classes to infer the full state of the system. On the other hand, MS-CGAN is an architecture derived from QST conditional-GAN (QST-CGAN)[39], which contain certain modifications from the original work to extend its reconstruction range to include mixed states. This network is improved for a better estimation of the density matrix of the system while extend reconstruction range to multiple type of quantum states. Both of these techniques can accurately construct many types of given simulated measurement and shown great promise in practical intermediate-scale quantum systems.
II Experimental data
@@
II-A Fock state
Eigenstates of the Fock basis:
@@ II-A Fock state
II-B Coherent state
Displaced vacuum states, characterized by the complex displacement amplitude $\alpha$:
@@ II-B Coherent state
II-C Thermal state
Mixed states where the photon number distribution follows super-Poissonian statistics:
@@ II-C Thermal state
II-D Cat state
Bosonic-code states consisting of superpositions of coherent states up to a normalization $N$ with projections $\Pi_{r}$ given by:
@@ II-E Num state
II-F Binomial state
Bosonic-code states constructed from a superposition of Fock states weighted by the binomial coefficients:
@@ II-F Binomial state
II-G GKP state
Finite Gottesmann-Kitaev-Preskill states, limits the lattice and adds a Gaussian envelope to make the state normalizable:
@@ II-G GKP state
III Architecture and training settings
III-A Restricted Feature Based Neural Network
@@
III-B Mixed State Conditional Generative Adversarial Network
Building upon the seminal work of QST-CGAN, we have developed a novel neural architecture for state prediction, utilizing the powerful and versatile TensorFlow [46] framework. Our proposed model is designed to accommodate a batch of measurements in the form of 32x32 density matrices, as well as label vectors representing the target state. To facilitate efficient processing, we have implemented a series of pre-processing steps, which include flattening the measurements to a 1024-length vector and one-hot encoding the label to a 7-length vector.
@@ IV Result
IV-A Restricted Feature Based Neural Network
@@
IV-B Mixed State Conditional Generative Adversarial Network
Figure 13: Training loss progression of MS-CGAN.
Figure 14: Validation fidelity progession of MS-CGAN during training process.
The MS-CGAN architecture exhibits a distinct characteristic when compared to the RFB-Net architecture, in that it is capable of converging to desired values within a training period consisting of no more than 100 iterations (Figure 13 and 14), despite requiring an extra label input in comparison to the RFB-Net architecture. Notably, the MS-CGAN architecture is primarily focused on reconstruction, but this does not detract from its robustness, given its swift convergence when trained on relevant data. It is therefore reasonable to conclude that the MS-CGAN architecture is a viable alternative to the RFB-Net architecture, given its impressive performance characteristics in terms of convergence speed and accuracy, despite having distinct differences in terms of input requirements and model focus.
IV-C Compare with existing methods
Figure 15: Validation fidelity progression of QST-CGAN on generated dataset.
Figure 16: Samples of QST-CGAN predictions.
@@ IV-C Compare with existing methods
In order to ensure a more robust and comprehensive comparison between our proposed model and the original model, we have undertaken a meticulous and rigorous experimental procedure that involves the use of a QST-CGAN model trained on the same dataset for a prolonged period of 100 epochs, utilizing the original configuration parameters and settings as described in the seminal paper. This approach is intended to provide a more nuanced and detailed analysis of the comparative performance of the two models, by subjecting both models to identical experimental conditions, and enabling us to assess their relative strengths and weaknesses in a more systematic and rigorous manner.
Upon conducting this extensive analysis, we have observed that the QST-CGAN model, while exhibiting notable strengths and advantages in certain areas, has encountered certain challenges and limitations in the reconstruction of specific batches of states, resulting in an unstable and erratic convergence on the dataset. These limitations are reflected in the average fidelity score, which is found to be hovering around a modest 0.19, indicating a suboptimal and unsatisfactory level of accuracy and precision. In contrast, our proposed solution has proven to be more superior and effective in overcoming these challenges, by demonstrating a higher degree of stability, robustness, and accuracy, as well as a more consistent and reliable performance.
IV-D Benchmarking with noise
The presence of noise, which refers to any unwanted fluctuations or disturbances in the experimental setup, is an inescapable factor that pervades most scientific investigations, and consequently poses significant challenges for accurate and reliable data analysis. Given this ubiquitous nature of noise, it is imperative that any method employed for state reconstruction is designed and optimized to be robust and resilient against various types of noise, to ensure that the final results are valid and trustworthy.
Moreover, it is essential to recognize that noise can infiltrate the system at multiple stages, leading to a potentially complex and nuanced interplay of factors that can affect the experimental outcome. Specifically, noise can manifest itself during the preparation of the state to be reconstructed, where errors in the preparation procedure can introduce small deviations in the actual state of interest, resulting in a distorted or altered version of the target state. Additionally, the measurement protocol itself can be a source of noise, as any calibration errors or other external influences may cause us to measure something other than what we intended to measure, leading to further inaccuracies in the final results. Finally, noise can arise during the data collection process itself, where amplification errors or photon shot noise can corrupt the signal, leading to further challenges in data interpretation and analysis. Therefore, a comprehensive and multifaceted approach is needed to account for the various sources of noise and ensure that any method for state reconstruction is adequately robust and resilient against these influences, to provide accurate and reliable results that can be used to advance scientific knowledge and understanding.
@@ IV-D Benchmarking with noise
Figure 17: Samples of noisy input and RFB-Net predictions.
@@
Figure 18: Samples of noisy input and MS-CGAN predictions.
In this research, our approach involves the utilization of various noise generating techniques, akin to the employed dataset, taken from QST-NN. In particular, we have employed three types of noise in our experiment, which include additive mixed state noise, photon loss noise, and pepper noise. Rather than introducing the noise during the training phase, whereby the model would gradually learn and adapt to it, we have opted to evaluate the performance of two distinct architectures devoid of exposure to noisy data. This is to ascertain whether our models can generalize well under these conditions or not.
Upon analyzing the sample input and prediction, it is evident that the RFB-Net exhibits a commendable performance in noisy environments in comparison to the MS-CGAN with regards to fidelity, albeit the reconstructed operator showcasing some dissimilarities from its original form. On the other hand, the MS-CGAN’s performance is more stable, particularly when dealing with pepper noise added samples, and shows superior feature extraction capabilities compared to RFB-Net.

### v3

@@ Abstract
Quantum state tomography is a crucial technique for characterizing the state of a quantum system, which is essential for many applications in quantum technologies. In recent years, there has been growing interest in leveraging neural networks to enhance the efficiency and accuracy of quantum state tomography. Still, many of them did not include mixed quantum state and pure states are arguably less common in practical situations. In this research paper, we present two neural networks based approach for both pure and mixed quantum state tomography: Restricted Feature Based Neural Network and Mixed States Neural Network, evaluate its effectiveness in comparison to existing neural based methods. We demonstrate that our proposed methods can achieve state-of-the-art results in reconstructing mixed quantum states from experimental data. Our work highlights the potential of neural networks in revolutionizing quantum state tomography and facilitating the development of quantum technologies. Source code is publicly available at https://github.com/luutn2002/uni-qst
@@ I Introduction
Quantum state tomography (QST) [15, 16, 17] aims to ascertain the unknown quantum state by conducting measurements on a finite set of identical copies of the system. If the state is characterized by the density matrix $\rho$, residing in a $d$-dimensional Hilbert space, approximately $O(d/\varepsilon)$ copies are necessary to achieve an estimate of $\rho$ with an error (measured as total variation distance) less than $\varepsilon$[18]. This underscores the considerable resource demands of QST for large-scale systems.
@@
In our research, we introduce QST with 2 type of neural networks (NNs) architecture: Restricted Feature Based Neural Network (RFB-Net) and Mixed States Neural Network (MS-NN). RFB-Net is a modified QST-NN[27] with added regression modules that can classify quantum states into specific specific class and pointing out its features of the quantum state within certain constraints, finally using these features and classes to infer the full state of the system. On the other hand, MS-NN is an architecture derived from QST conditional-GAN’s generator (QST-CGAN)[28], which contain certain modifications from the original work to extend its reconstruction range to include mixed states. This network is improved for a better estimation of the density matrix of the system while extend reconstruction range to multiple type of quantum states. Both of these techniques can accurately construct many types of given simulated measurement and shown great promise in practical intermediate-scale quantum systems.
II Background
In this section we will discuss some background knowledge related to our research, including knowledge related to QST and usage of neural network (NN) methods in QST.
II-A Quantum state tomography
Quantum state tomography (QST) is a fundamental technique in quantum information science, providing a method to reconstruct the quantum state of a system based on measurement outcomes. The process of QST involves performing a series of measurements on identically prepared quantum systems and then using statistical methods to infer the quantum state, typically represented as a density matrix $\rho$.
In the context of QST, the primary targets for reconstruction are typically either pure or mixed quantum states. For a pure quantum state, its density matrix is Hermitian, positive semi-definite, has a trace of one, and can be expressed as the outer product of a state vector with itself:
$\rho=|\psi\rangle\langle\psi|$ (1)
Conversely, if the quantum state is mixed, it is represented as:
$\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ (2)
where $p_{i}$ denotes the probability of the system being in the pure state $|\psi_{i}\rangle$. Since mixed states cannot be described by a single ket vector[29], converting quantum states to density matrices allows for a unified approach to their reconstruction.
During the reconstruction process of pure quantum states, it is common to simulate a Cholesky decomposition [27, 28] of the density matrix:
$\rho=LL^{\top}$ (3)
where $L$ represents a lower triangular matrix with complex entries, while the diagonal elements are real. This structure guarantees that the resulting matrix maintains Hermiticity and positivity.
One of the common method to obtain the density matrix $\rho$ for QST process is to perform a single-shot measurement on positive operator-valued measures (POVMs) $\{\mathcal{O}_{i}\}$. The frequencies $d_{i}$ of various measurement outcomes are proportional to the expectation value $\text{tr}(O_{i}\rho)$ and constitute our data $d$. Hence, the reconstruction problem can be formulated as an inversion problem [30, 31, 32]:
$d=A\rho_{f}$ (4)
where $A$ represents the sensing matrix determined by the choice of measurement operators, and $\rho_{f}$ is the vectorized density matrix. The ability to invert Equation 4 relies on the selection of measurement operators. A set of measurement operators that allows inversion and thus complete characterization of the state is termed informationally complete (IC) [33]. For a state in an $N$-dimensional Hilbert space, approximately $N^{2}$ number of POVMs might be necessary for IC, with each measurement might needs multiple repetitions to accumulate sufficient statistics. However, with prior knowledge about the state (such as $\rho$ is low rank or or contain certain zero elements), the number of required measurements can be reduced through judicious selection.
The process of reconstructing $\rho$ from d is an estimation challenge that can be approached through various methods. Common techniques include linear inversion [34], maximum-likelihood estimation (MLE)[35, 36] and Bayesian methods[37, 38]. While linear inversion is straightforward, it may fail due to data noise or a high condition number of sensing matrix $A$ [32], potentially leading to unphysical results such as negative diagonal elements in the density matrix. Thus, statistical methods like MLE or Bayesian estimation are preferred. These methods estimate $\rho^{\prime}$ by maximizing the likelihood function:
$L(\rho^{\prime}\mid\textbf{d})=\prod_{i}[\text{tr}(\rho^{\prime}O_{i})]^{d_{i}}$ (5)
For continuous-variable outputs, where $d_{i}$ is a real number, appropriate binning is required for MLE application [39]. Alternatively, minimizing the mean squared error between the observed and expected values is another approach[40].
Though MLE ensures a physical $\rho$, it does not provide error bars to measure uncertainty. Recent discussions suggest that MLE may not be optimal and could be inadmissible for metrics like fidelity, mean-squared error, and relative entropy [41]. In contrast, Bayesian methods can quantify uncertainty by using a prior probability distribution $\pi(\rho)$ [37, 38]. The prior $\pi_{0}(\rho)$ is typically chosen to be uniform or as uninformative as possible, and it is updated using Bayes’ theorem with the likelihood $L(\rho^{\prime}\mid d)$ to yield a posterior distribution $\pi_{f}(\rho)\propto L(\rho\mid d)\pi_{0}(\rho)$. The best estimate of the state is then the mean over all states $\rho_{\mu}$, weighted by the posterior distribution:
[…]
Other methods for optimizing the likelihood function and estimating the density matrix include diluted MLE[42], compressed sensing (CS)[43, 44], and projected gradient descent (PGD)[45, 46]. CS methods are motivated by parameter-counting arguments, suggesting that only $O(rN)$ measurements are needed, where $r$ is the rank of the density matrix [47]. For low-rank states, such as pure states affected by local noise, recent CS modifications for adaptive tomography [44, 48] require only the dimension of the density matrix, which is an improvement over traditional CS that requires an estimate of $r$.
Effective reconstruction can also be achieved by using models or ansätze that describe quantum states with fewer parameters than the general $N^{2}-1$. Techniques like matrix-product-state (MPS) [49, 50] and tensor-network (TN) tomography [51] provide efficient models for states, and permutationally invariant tomography [52] utilizes symmetries of the density matrix for mixed state reconstruction. Other improvements assume specific noise models, such as additive Gaussian noise[40], but these methods often lack general applicability.
An alternative approach is projected gradient descent (PGD)[45, 46], where a cost function differentiates between model-predicted and true data to apply gradient-based optimization. PGD is advantageous because it converges quickly to the MLE state in a broader range of scenarios, even for ill-conditioned problems.
II-B Neural network based approach in quantum state tomography
Neural networks have recently gained prominence as a robust method for extracting compact representations from high-dimensional datasets [53, 54, 55]. In the realm of experimental quantum science, these methods have been successfully employed for tasks such as classifying experimental snapshots [56, 57] and qubit readout [58]. This data-driven methodology can also be extended to tomographic tasks and are showing promising results in QST. These methods either use neural networks as ansatz to predict measurement probabilities [26, 52, 59] or directly estimate the density matrix $\rho$ [60]. Recent theoretical advances have shown that a generative model, specifically the restricted Boltzmann machine (RBM), can accurately reconstruct quantum states and observables using synthetic datasets produced by numerical algorithms [61].
Noticeably, [62, 63] had explored the possibility of applying RBM for mixed state tomography. To reconstruct a mixed state density matrix $\rho_{S}$, the approach mirrors method used for pure states. We construct a neural network with parameters $\theta$, and for a fixed basis $|v\rangle$, the density operator is represented by the matrix elements $\rho(\theta,v,v^{\prime})$, which are determined by the neural network. Consequently, the mapping from a neural network with parameters $\theta$ to a density operator can be expressed as:
$\rho(\theta)=\sum_{v,v^{\prime}}|v\rangle\rho(\theta,v,v^{\prime})\langle v^{\prime}|$ (7)
To achieve this, we employ the purification method for density operators. In this framework, the environment is modeled by additional hidden neurons $e_{1},\ldots,e_{m}$, in addition to the hidden neurons $h_{1},\ldots,h_{l}$. The purification $|\Psi_{SE}\rangle$ of $\rho_{S}$ is represented by the parameters of the neural network, which we still denote by $\theta$, such that:
$|\Psi_{SE}\rangle=\sum_{v}\sum_{e}\Psi_{SE}(\theta,v,e)|v\rangle|e\rangle$ (8)
By tracing out the environmental degrees of freedom, the density operator is determined by the network parameters as:
$\rho_{S}=\sum_{v,v^{\prime}}\left(\sum_{e}\Psi_{SE}(\theta,v,e)\Psi^{*}_{SE}(\theta,v^{\prime},e)\right)|v\rangle\langle v^{\prime}|$ (9)
Despite these advances, a versatile solution for large-scale NN-based QST solution that can accommodate various reconstruction scenarios has not yet been thoroughly explored. We aim to address this gap in the following sections.
III Data and noise model
In this section, we will discuss the employed quantum states used to gather measurements for our reconstruction experiments and introduced noise during benchmarking process.
III-A Experimental data
@@
III-A1 Fock state
Eigenstates of the Fock basis
@@ III-A1 Fock state
III-A2 Coherent state
Displaced vacuum states, characterized by the complex displacement amplitude $\alpha$
@@ III-A2 Coherent state
III-A3 Thermal state
Mixed states where the photon number distribution follows super-Poissonian statistics
@@ III-A3 Thermal state
III-A4 Cat state
Bosonic-code states consisting of superpositions of coherent states up to a normalization $N$ with projections $\Pi_{r}$ given by
@@ III-A5 Num state
III-A6 Binomial state
Bosonic-code states constructed from a superposition of Fock states weighted by the binomial coefficients
@@ III-A6 Binomial state
III-A7 GKP state
Finite Gottesmann-Kitaev-Preskill states, limits the lattice and adds a Gaussian envelope to make the state normalizable
@@ III-A7 GKP state
III-B Experimented noise models
Noise is an unavoidable element in most experimental setups. Consequently, techniques for reconstruction need to be sufficiently resilient to different forms of noise. In this subsection, we delineate the various types of noise employed to evaluate our neural network-based reconstruction methods.
Noise can affect the process at multiple stages. Initially, errors during the preparation of the state to be reconstructed can result in a slightly altered state, $\rho\rightarrow\rho_{\text{noisy}}$ (state-preparation errors). Secondly, errors in the measurement protocol, such as calibration inaccuracies, may cause deviations in the measurements, $\{\mathcal{O}_{i}\}\rightarrow\{\mathcal{O}_{\text{noisy},i}\}$ (measurement errors). Finally, errors may arise during data collection, such as those introduced by signal amplification or photon shot noise, leading to corrupted data, $\mathbf{d}\rightarrow\mathbf{d}_{\text{noisy}}$ (data errors).
State-preparation and measurement (SPAM) errors can be systematic and thus challenging to correct. Recently, deep neural networks have been shown to be effective in learning and correcting such errors [69] by training a supervised model to correct the data, $\mathbf{d}_{\text{noisy}}\rightarrow\text{DNN}\rightarrow\mathbf{d}$. In this context, neural networks act as sophisticated filters that denoise experimental data while remaining agnostic to the underlying SPAM noise. In this paper, for reconstruction, we demonstrate that the neural network approach is robust against the various types of SPAM and data errors described below.
III-B1 Mixed state noise
In many experimental settings, quantum states are influenced by thermal and other types of environmental noise. To account for this, we represent the noise by using mixed states:
$\rho_{\text{mixed}}=(1-\zeta)\rho+\zeta\rho_{\text{random}},$ (16)
where $\zeta\in[0,0.5]$. For the classification task, the correct label of a mixed state is determined by the class to which $\rho$ belongs. In the reconstruction task, the goal is not to reconstruct $\rho$ itself, but rather $\rho_{\text{mixed}}$, as this reflects the actual state produced during the experiment.
III-B2 Photon loss noise
When an optical quantum state is generated in a lossy resonator, photons can escape from the resonator before the state measurement is completed. We model this photon loss by allowing the initial state to evolve over a time $\tau$ according to the master equation:
$\dot{\rho}=-\frac{i}{\hbar}[H,\rho]+\gamma\mathcal{L}[a]\rho$ (17)
where $H=\hbar\omega a^{\dagger}a$ is the free resonator Hamiltonian, $\omega$ is the resonator frequency, $\gamma$ denotes the photon loss rate, and $\mathcal{L}[a]\rho=a\rho a^{\dagger}-\frac{1}{2}a^{\dagger}a\rho-\frac{1}{2}\rho a^{\dagger}a$. Similar to the mixed states discussed, in the classification task, the correct label corresponds to the class of $\rho(t=0)$. In the reconstruction task, however, this noise is not necessarily considered an error, as the objective is to reconstruct $\rho(t=\tau)$.
III-B3 Pepper noise
Salt-and-pepper noise is characterized by sudden, drastic changes in the signal at a limited number of points. To model dead pixels or missing data, we use pepper noise, where a random fraction of data points is selected and set to zero.
IV Architecture and training settings
In this section, we will introduce the architectural design of NN-based models and some notable training settings used in our experiments.
IV-A Restricted Feature Based Neural Network
@@
IV-B Mixed State Neural Network
Building upon the seminal work of QST-CGAN[28], we derived MS-NN from QST-CGAN’s generator[28], utilizing the powerful and versatile TensorFlow [71] framework. Our proposed model is designed to accommodate a batch of measurements in the form of 32x32 density matrices, as well as label vectors representing the target state. To facilitate efficient processing, we have implemented a series of pre-processing steps, which include flattening the measurements to a 1024-length vector and one-hot encoding the label to a 7-length vector.
@@ V Result
V-A Restricted Feature Based Neural Network
@@
V-B Mixed State Neural Network
Figure 13: Training loss progression of MS-NN.
Figure 14: Training fidelity progression of MS-NN during training process.
The MS-NN architecture exhibits a distinct characteristic when compared to the RFB-Net architecture, in that it is capable of converging to desired values within a training period consisting of no more than 100 iterations (Figure 13 and 14), despite requiring an extra label input in comparison to the RFB-Net architecture. Notably, the MS-NN architecture is primarily focused on reconstruction, but this does not detract from its robustness, given its swift convergence when trained on relevant data. It is therefore reasonable to conclude that the MS-NN architecture is a viable alternative to the RFB-Net architecture, given its impressive performance characteristics in terms of convergence speed and accuracy, despite having distinct differences in terms of input requirements and model focus.
V-C Compare with existing methods
Figure 15: Training fidelity progression of QST-CGAN[28] when training on generated dataset.
Figure 16: Samples of QST-CGAN[28] predictions.
@@ V-C Compare with existing methods
In order to ensure a more robust and comprehensive comparison between our proposed model and the original model, we have undertaken a meticulous and rigorous experimental procedure that involves the use of a QST-CGAN[28] model trained on the same dataset for a prolonged period of 100 epochs, utilizing the original configuration parameters and settings as described in the seminal paper. This approach is intended to provide a more nuanced and detailed analysis of the comparative performance of the two models, by subjecting both models to identical experimental conditions, and enabling us to assess their relative strengths and weaknesses in a more systematic and rigorous manner.
Upon conducting this extensive analysis, we have observed that the QST-CGAN[28] model, while exhibiting notable strengths and advantages in certain areas, has encountered certain challenges and limitations in the reconstruction of specific batches of states, resulting in an unstable and erratic convergence on the dataset. These limitations are reflected in the average fidelity score, which is found to be hovering around a modest 0.19, indicating a suboptimal and unsatisfactory level of accuracy and precision. In contrast, our proposed solution has proven to be more superior and effective in overcoming these challenges, by demonstrating a higher degree of stability, robustness, and accuracy, as well as a more consistent and reliable performance.
V-D Benchmarking with noise
Rather than introducing the noise during the training phase, whereby the model would gradually learn and adapt to it, we have opted to evaluate the performance of two distinct architectures devoid of exposure to noisy data. This is to ascertain whether our models can generalize well under these conditions or not.
@@ V-D Benchmarking with noise
Figure 17: Samples of noisy input measurements and RFB-Net generated state measurements.
@@
Figure 18: Samples of noisy input measurements and MS-NN generated state measurements.
Upon analyzing the sample input and prediction, it is evident that the RFB-Net exhibits a commendable performance in noisy environments in comparison to the MS-NN with regards to fidelity, albeit the reconstructed operator showcasing some dissimilarities from its original form. On the other hand, the MS-NN’s performance is more stable, particularly when dealing with pepper noise added samples, and shows superior feature extraction capabilities compared to RFB-Net.